OpenStudy (loser66):

How to simplify \((2\overline{z}-i)(2z+i) =|2z+i|^2\) Please, help

OpenStudy (loser66):

nvm got it

OpenStudy (anonymous):

let \[z=x+i y, \space \overline z=x-iy\] \[\overline z \times z=x^2+y^2\] left hand side: \[(2\overline{z}-i)(2z+i)=[4(x^2+y^2)+2i(x-iy)-2i(x+iy)+1]\] \[[4(x^2+y^2)+2i(x-iy)-2i(x+iy)+1]=[(4x^2+4y^2)+2ix+2y-2ix+2y+1]\] \[[4x^2+4y^2)+2ix+2y-2ix+2y+1]=[(4x^2+4y^2)+4y+1]\] right hand side: \[|2z+i|^2=(\sqrt{(2x)^2+i(2y+1)^2}=4x^2+4y^2+4y+1\] so LHS=RHS